OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

first you want to get the absolute value by itself

OpenStudy (anonymous):

@kaylamarie081

OpenStudy (anonymous):

I just really need quick answers, Im behind and at risk of being withdrawn :(

OpenStudy (anonymous):

|x+5|=4 x+5=4 or x+5=-4 x=-1 or x=-9

OpenStudy (anonymous):

Can you help with a few more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

] Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (anonymous):

the question mark things are all negative signs

OpenStudy (anonymous):

|2x+6|=24 2x+6=24 or 2x+6=-24 2x=18 or 2x=-30 x=9 or x=-15

OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution